# Conceptualization, Data Curation, Software, Visualization, Writing - original draft

### 01 Mar 2026 v5

@@ 2.1 Loss of tensile strength (LoTS) for ACSR lines
The accumulated LoTS is obtained by integrating temperature exposure over time, enabling direct quantification of ETO-induced degradation. Figure 1 illustrates the calculation process using temperature–time exposure pairs ordered from low to high temperature. At each step, prior exposures are converted into an equivalent time at the current (higher) temperature, added to the new exposure, and the remaining tensile strength is updated via the LoTS model. Repeating this process yields the total LoTS at the highest temperature.
2.2 Probability of Failure from LoTS
The consequence of ETO can be modeled through the increased probability of failure (PoF) of the Overhead conductor. The instantaneous a posteriori PoF was used, which reflects the PoF given the survival until the current conductor LoTS $L^{\text{ACSR}}$:
@@ 2.2 Probability of Failure from LoTS
where $f(\cdot)$ is the probability density function (PDF) for modeling the failure. The use of the a posteriori PoF enables modeling of failure probability considering the events leading up to $L^{\text{ACSR}}$ Teh2017.
2.3 Piecewise depreciation cost for overhead lines
The increased risk due to ETO can be translated into economic depreciation. Utilities choose to replace conductors when the risk of the conductor failure is larger than the reconductoring costs AusNet2025.
Thus, the replace timing of the conductor $L^{\text{ACSR, End}}_{l}$ can be balanced using the following equation:
@@ 2.3 Piecewise depreciation cost for overhead lines
where $C^{\text{Fail}}_{l}$ is the cost of failure of line $l$ and $C^{\text{Replace}}_{l}=B_{l}\cdot S_{l}\cdot D_{l}$. Here, $B_{l}$ is the initial cost factor of line $l$ (\$/MVA-km), $S_{l}$ its capacity (MVA), and $D_{l}$ its length (km). Subscript $t$ denotes the time of interest. Since the typical end-of-life value of LoTS lines are given as 10% ($L^{\text{ACSR, End}}_{l}=10$) Adomah2000, miso2021workshop, equation (4) can be expressed as follows:
@@
Then the cost of conductor depreciation from one hour ETO (or the risk due to the incremental LoTS: $L^{\text{ACSR}}_{l,t}(\tau)-L^{\text{ACSR}}_{l,t-1}$) can be modeled by subtracting equation 5 for consecutive hours:
$C^{\text{Dep}}_{l,t}(\tau)=C^{\text{Replace}}_{l}\cdot\frac{\pi(L^{\text{ACSR}}_{l,t}(\tau))-\pi(L^{\text{ACSR}}_{l,t-1})}{\pi(10)}$ (6)
subscript $t$ denotes the time of interest. This formulation implies that $C^{\text{Dep}}_{t}(\tau)$ varies with the state of the conductor as $L^{\text{ACSR}}_{t}(\cdot)$ is state-dependent. While the depreciation model in ((6)) can be applied in post-hoc evaluation, this research aims to integrate it into the UC problem with DLR which involves the real-time decision-making. Thus, we approximate the expected cost of a one-hour ETO by formulating the 1-hour ETO costs at different temperatures as a piecewise function prior to solving the daily CHA-UC problem. This function is convex, as higher temperature yield steeper depreciation slopes, and is calculated separately for each conductor. Figure 2 depicts the depreciation cost function and its piecewise counterpart for an old transmission line with many ETO histories (old) and a transmission line with less ETO histories (new). The old line exhibits lower depreciation costs because the cost becomes less sensitive to additional ETO exposure as LoTS accumulates.
@@
Figure 3: Illustration of a linear fit to the current-conductor temperature relationship. The temperature from the IEEE HBE and the linear fit assuming ACSR Finch, a perpendicular wind speed of 2.7 m/s and ambient temperature of 9.4°C is shown. Conductor temperature limit is set as 95 °C; in practice, it varies depending on the transmission system owner Brown2024.
2.4 Heat balance equation
The steady-state heat balance equation (HBE) IEEE738 is used to obtain conductor temperature $\tau$. Since our focus is on hour-scale unit commitment, transient effects are neglected: conductor reaches 95% of the steady-state temperature within approximately 42 minutes Park2018. The steady-state HBE is expressed as:
@@ 2.4 Heat balance equation
where $q_{c}$, $q_{r}$, and $q_{s}$ denote convective and radiative heat losses and solar heat gain, respectively. $I^{2}R(\tau_{\text{avg}})$ is the Joule heating term. Among these, wind speed has the strongest influence on the DLR, followed by ambient temperature and solar radiation Wallnerstrom2015. Additional information on the details of the HBE can be found in IEEE738.
2.5 Linear temperature approximation
@@
where $p_{l,t}$ is the line power flow. Since voltage is assumed constant in the DC power flow, $A_{l,t}$ is scaled by $\frac{1}{\sqrt{3}V^{\text{Nom}}_{l}}$ from the current–temperature fit. The two inequalities account for bidirectional flow, ensuring $\tau_{l,t}$ is an upper bound on the actual conductor temperature. The use of a linear approximation makes it unnecessary to explicitly express the absolute value of current in optimization, simplifying the conductor temperature model.
3 Conductor health-aware unit commitment
Adopting DLR pushes line temperatures toward thermal limits, amplifying ETO risk in the presence of forecast errors. We propose the CHA-UC framework, which minimizes the sum of expected conductor depreciation and system operating costs to balance reliability and economics. This enables the system operator to provide the most effective commitment decisions and dispatch, using DLR.
3.1 Baseline UC model
Before presenting the CHA-UC framework, the deterministic baseline model for DLR and SLR is introduced. Day-ahead UC and real-time operation is reflected in the model. The UC problem fixes commitment decisions day before the actual operation, using day-ahead forecasts and considering physical constraints of generators. The enforcement of UC constraints lead to a more comprehensive benchmark of results Park2018. The operation step adjusts the generation of available generators under ramping constraints and fixed commitments to account forecast error.
@@ 3.1 Baseline UC model
which is the sum of production cost, startup cost and the shut down cost. $\Xi:=\{\mathbf{g,p,u,}\,\boldsymbol{\theta,\delta}\}$
The day-ahead UC constraints, hereafter referred to as the UC constraints, are adapted from the formulation of Morales-España et al. Morales2013:
@@
$\begin{split}&\sum_{l\mid s(l)=b}\hskip-4.2679ptp_{l,t}-\hskip-4.2679pt\sum_{l\mid r(l)=b}\hskip-4.2679ptp_{l,t}+\sum_{i\in\mathcal{I}_{b}}g_{i,t}+\hskip-4.2679pt\sum_{w\in\mathcal{W}_{b}}(\hat{P}^{\rm wind}_{w,t}-p^{\text{cur}}_{w,t})\\ &=P^{\text{d}}_{b,t}-p^{\text{shed}}_{b,t},\quad\forall b\in\mathcal{B},t\in\mathcal{T},\end{split}$ (11b)
@@
Once the day-ahead dispatch is fixed, the real-time operation minimizes the additional cost of balancing actual wind generation and DLR realizations:
@@
where $\Xi:=\{\mathbf{g}^{\mathrm{adj}},\mathbf{p}^{\mathrm{adj}},\mathbf{r}^{\mathrm{+,-}},\boldsymbol{\theta}^{\mathrm{adj}}\}$. Reserve activations are penalized through $C_{i}^{\mathrm{r}\pm}$, and the quadratic term captures nonlinear production costs. This objective represents the operator’s total adjustment cost in real time. The constraints of the operation stage are as follows:
@@
3.2 Conductor health-aware unit commitment
Given the linear fit and the piecewise depreciation cost function in Section 2, the objective function of the CHA-UC day-ahead problem is as follows:
[…]
The objective includes the operation cost from (9) and the scenario stochastic costs: the expected conductor depreciation cost, expected load shed costs and the expected reserve activation costs. The operation cost and the expected conductor depreciation cost use same units (\$), since the conductor depreciation model is designed to be incorporated to the UC problem. This eliminates the need for any additional scaling or conversion coefficient, which simplifies the optimization and improves interpretability. The decision variables are identical to the baseline. The additional term in (14) represents the expected values from considering stochasticity. The sample average approximation was used to calculate the expected value. $C^{\text{Dep}}_{l,t}$ is a piecewise function of the depreciation cost of ETO in Section 2.3, implemented using SOS2 constraints.
@@ 3.2 Conductor health-aware unit commitment
$\begin{split}&\sum_{l\mid s(l)=b}\hskip-4.2679pt\hat{p}^{\text{adj}}_{l,t,s}-\sum_{l\mid r(l)=b}\hskip-4.2679pt\hat{p}^{\text{adj}}_{l,t,s}+\sum_{i\in\mathcal{I}_{b}}\hat{g}^{\text{adj}}_{i,t,s}+\hskip-4.2679pt\sum_{w\in\mathcal{W}_{b}}\hskip-4.2679pt(\hat{P}^{\rm wind}_{w,t}\\ &-\hat{p}^{\text{cur,adj}}_{w,t}-\omega_{w,t,s})=P^{\text{d}}_{b,t}-\hat{p}^{\text{shed,adj}}_{b,t,s},\quad\forall b\in\mathcal{B},t\in\mathcal{T},\end{split}\hskip-8.53581pt$ (16f)
@@
The surrogate of the “wait-and-see” operation stage is formed to reflect the stochasticity in the “here-and-now” day-ahead UC. The surrogate corresponding (13a)–(13g) are (16a)–(16g), respectively. They all represent the same operational constraints, but the latter are not actual; they represent the constraints that are applied to scenarios. Thus, in (16f) and (16g), $\omega_{w,t,s}$ represents the wind forecast error for each scenario. Equation (16h) and (16i) constrain $\tau_{l,t,s}$ using the linear conductor-temperature approximation, as formulated in (8). $\tau_{l,t,s}$ will be fixed to the lower bound as $C^{\text{Dep}}_{l,t}$ is a convex function, working as a proxy of conductor temperature. $\xi_{l,t,s}$ is the error of the DLR forecast, but instead of subtracting it from the DLR, we add it to the line flow.
@@
The additional conductor temperature constraints enable the conductor temperature to be reflected in the objective function. Analogous to equations (16h)-(16i), equations (19a)-(19b) enforce $\tau_{l,t}$ to reflect the conductor temperature before the operation stage.
3.3 CHA-UC Timeline
Figure 4: Schematic of the CHA-UC optimization timeline. Processes that require actions outside optimization is colored with gray. Equations and factors are marked with its references.
The CHA-UC timeline is shown in Figure 4. A UC problem is solved 18:00 the day before, and an operation problem is solved real-time (or hour-ahead). The loading-temperature linear fit should be updated everyday before the UC problem is solved to reflect the most recent conductor condition based on the latest forecasts. After operation, the depreciation cost function is updated for the next optimization period.
4 Case Study
Two case studies are performed on the synthetic Texas 123-bus backbone transmission system (TX-123BT) Lu2025 using 2021-2022 weather data. Case Study I runs an annual simulation with uncertain DLR forecasts (with errors) and perfect wind power (WP) forecasts (no errors) to illustrate CHA-UC. Case Study II evaluates CHA-UC in a more realistic setting with uncertainty in both WP and DLR forecasts to analyze their correlation effects.
Table 1: Summarized results of Case Study I and Case Study II. The best result for each cost is in bold.
@@ 4 Case Study
Spring CHA 14.29 0.09 0.23 11.7 14.60 15.29 1.72 0.77 0.49 116.8 18.26
DLR 14.17 1.31 10.3 15.47 13.73 2.71 3.98 11.67 77.2 32.09
SLR 14.57 0.12 108.4 14.69 14.15 2.29 2.85 8.88 92.8 28.17
QRF 14.25 0.69 29.9 14.94 13.90 2.55 3.41 8.32 82.5 28.17
Summer CHA 29.99 0.12 0.13 11.9 30.21 30.79 2.56 1.39 0.14 52.7 34.87
DLR 29.63 0.95 3.4 30.57 28.48 4.12 5.15 10.81 28.9 48.56
SLR 29.96 0.46 64.8 30.42 28.84 3.68 4.28 8.35 38.9 45.14
QRF 30.35 0.19 32.94 30.55 29.63 3.77 4.61 6.32 34.9 44.33
Fall CHA 22.08 0.07 0.01 24.8 22.15 22.83 2.27 0.97 0.01 69.2 26.09
DLR 21.85 0.49 18.3 22.34 20.89 3.65 3.43 12.68 47.7 40.65
SLR 22.26 0.04 108.2 22.30 21.28 3.26 2.78 7.27 77.5 34.60
QRF 22.01 0.07 44.7 22.08 21.16 3.48 3.06 7.02 51.9 34.71
Winter CHA 18.50 0.05 0.13 1.1 18.67 19.28 1.91 0.61 0.00 75.6 21.81
DLR 18.39 0.65 4.4 19.04 17.72 3.01 3.02 8.70 48.6 32.44
SLR 18.86 0.00 50.9 18.86 18.15 2.61 2.06 4.98 57.0 27.80
QRF 18.44 0.27 4.3 18.71 17.81 2.94 2.81 6.01 40.8 29.56
Annual CHA 21.23 0.08 0.12 12.5 21.44 22.07 2.12 0.94 0.16 78.6 25.29
DLR 21.04 0.85 9.1 21.89 20.23 3.38 3.90 10.98 50.6 38.49
SLR 21.44 0.16 83.4 21.60 20.63 2.96 3.00 7.39 66.6 33.99
QRF 21.29 0.31 28.2 21.60 20.65 3.19 3.48 6.92 52.7 34.24
4.1 Preliminaries
4.1.1 Simulation Settings
In all case studies, a set of 21 lines was selected for DLR application [^1] Lines were screened by solving an annual DCOPF with SLR and selecting those with the highest dual values of the line-flow constraint; lines longer than 100 km were excluded because their limits are dominated by stability.. All other lines used SLR limits. Each 345 kV line was modeled as ACSR Finch, and ampacity was adjusted to fit the TX-123BT system through a line multiplier. All optimization models were implemented in Julia/JuMP Lubin2023 and solved with Gurobi. The LoTS limit was set to 10% (Section 2.3). The lifetime cost factor $B_{l}$ was based on a 3190 MVA, 400 kV AC line with a cost of 750 GBP/MVA-km Parsons2012. After excluding tower and loss components, the adjusted cost factor ($B_{l}$) was \$936/MVA-km. The PDF modeling line failure $f(\cdot)$ was modeled as a normal distribution following $\mu=0.1,\sigma=0.05$. The loading-temperature linear fit was sampled from 95 °C and 150 °C, and the conductor depreciation piecewise function was sampled every 10 °C from 95 °C. Conductor health was initialized to fit a mean LoTS rate of 1% (a sensible rate from the results of Adomah2000) by creating synthetic historical ETO exposures from a normal distribution. The ETO exposure data is passed on between different horizons to update the conductor state properly. For DLR security margin, line limits were constrained as $K_{l,t}\hat{\overline{F}}^{\mathrm{DLR}}_{l,t}=\max(0.8\hat{\overline{F}}^{\mathrm{DLR}}_{l,t},\,\underline{F}^{\mathrm{SLR}}_{l})$. The load profile of 2021 given in the TX-123BT test system was used, assuming no uncertainties in the load forecast. The value of lost load, $C^{\text{VoLL}}$, is assumed to be \$3500/MWh following MISO before 2025/05 Kirschen2018. Reserve activation costs followed Dvorkin2025: $C_{i}^{\mathrm{r}+}$ equals three times, and $C_{i}^{\mathrm{r}-}$ equals 0.5 times, the linear cost coefficient of generator $i$.
4.1.2 Data Collection
Weather data for 2021–2022 were obtained from the NOAA’s high resolution rapid refresh model Benjamin2016, which provides 3 km spatial resolution and 1 h temporal resolution. Forecasts issued at 18:00 the previous day were used as day-ahead inputs, while analysis data represented actual operation. Weather data from 2022 were used to generate stochastic scenarios, and 2021 data were used for annual simulations. WP profiles were derived from wind speeds using the standard power curve Giorsetto1983, adjusted to hub height via roughness-length correction. Uniform turbine specifications (cut-in and cut-out speeds) were assumed. For each transmission line, the hourly minimum DLR across all 3 km $\times$ 3 km grid cells intersecting the line corridor was used as the representative value. The parameters in (7) followed IEEE IEEE738, with solar heat gain fixed at the worst-case condition for Texas (latitude 31°, 1000 W/m2, day 180, and maximum projected area). The conductor thermal limit was set to 95°C.
Figure 5: Histograms comparing the distribution of actual and scenario errors. (a) DLR forecast error, (b) WP forecast error
4.1.3 Scenario Generation for CHA-UC
To generate representative forecast-error scenarios for both WP and DLR while preserving their correlation, we construct a unified trajectory dataset where each record corresponds to one historical day and contains the 24-hour forecast errors. To ensure tail coverage, several historically extreme trajectories (large positive and negative errors) are first seeded. A greedy max–min selection then iteratively adds the trajectory with the largest minimum distance to the existing set, maximizing overall diversity. Finally, we assign probabilities via the nearest-neighbor weighting: each historical trajectory is mapped to its closest selected scenario, and each scenario’s probability equals the fraction of histories in its cell. This yields a scenario set that reflects the empirical distribution of forecast errors.
A total of 20 scenarios were generated from the historical data. Figure 5 shows that the resulting distribution slightly overestimates the empirical tails, ensuring adequate tail-risk representation. The correlation structure was also well preserved–the average empirical correlation was 0.04, compared to 0.05 across the generated scenarios–demonstrating effective reproduction of inter-asset dependencies.

### 14 Apr 2026 v6

@@ II-A Loss of tensile strength (LoTS) for ACSR lines
The accumulated LoTS is obtained by integrating temperature exposure over time, enabling direct quantification of ETO-induced degradation.
II-B Probability of Failure from LoTS
The consequence of ETO can be modeled through the increased probability of failure (PoF) of the Overhead conductor. Instantaneous a posteriori PoF is used, reflecting the PoF given the survival until conductor LoTS $L^{\text{ACSR}}$:
@@ II-B Probability of Failure from LoTS
where $f(\cdot)$ is the probability density function (PDF) for modeling the failure. The use of (3) enables modeling failure probability considering the events leading up to $L^{\text{ACSR}}$ [Teh2017].
II-C Piecewise depreciation cost for overhead lines
The increased risk due to ETO is translated into economic depreciation. Utilities replace conductors when failure risk exceeds reconductoring cost [AusNet2025], yielding:
@@ II-C Piecewise depreciation cost for overhead lines
where $C^{\text{Fail}}_{l}$ is the cost of failure of line $l$ and $C^{\text{Replace}}_{l}=B_{l}\cdot S_{l}\cdot D_{l}$. Here, $B_{l}$ is the cost factor of $l$ (\$/MVA-km), $S_{l}$ its capacity (MVA), and $D_{l}$ its length (km). Using the typical end-of-life threshold $L^{\text{ACSR, End}}_{l}=10$ [Adomah2000, miso2021workshop], this becomes
@@
By subtracting equation (5) for a certain timestep, the depreciation (or risk) cost due to incremental LoTS is then:
$C^{\text{Dep}}_{l,t}(\tau)=C^{\text{Replace}}_{l}\cdot\frac{\pi(L^{\text{ACSR}}_{l,t}(\tau))-\pi(L^{\text{ACSR}}_{l,t-1})}{\pi(10)},$ (6)
where subscript $t$ denotes the time of interest. Because $C^{\text{Dep}}_{t}(\tau)$ depends on the conductor state, its direct use is limited to post-hoc evaluation. To embed depreciation costs into the daily UC problem, this study constructs conductor-specific convex piecewise approximations of the one-hour ETO cost as a function of temperature. Figure 1 compares the exact and piecewise depreciation costs for two lines; the old line exhibits lower incremental costs since conductor health becomes less sensitive to ETO exposure as LoTS accumulates.
@@
II-D Heat balance equation
The steady-state heat balance equation (HBE) [IEEE738] is used to obtain conductor temperature $\tau$. Since our focus is on hour-scale unit commitment, transients are neglected: conductor reaches 95% of the steady-state temperature within approximately 42 minutes [Park2018]. The steady-state HBE is:
@@ II-D Heat balance equation
where $q_{c}$, $q_{r}$, and $q_{s}$ denote convective and radiative heat losses and solar heat gain, respectively. $I^{2}R(\tau_{\text{avg}})$ is the Joule heating term. Among these, wind speed (convection) has the strongest influence on the DLR, followed by ambient temperature [Wallnerstrom2015].
II-E Linear temperature approximation
@@
Figure 2: Illustration of a linear fit to the current-conductor temperature relationship, given a perpendicular wind speed of 2.7 m/s and ambient temperature of 9.4°C. Conductor temperature limit is set as 95 °C; in practice, it depends on the transmission owner [Brown2024].
@@
where $p_{l,t}$ is the line power flow. Since voltage is assumed constant in the DC power flow, $A_{l,t}$ is scaled by $\frac{1}{\sqrt{3}V^{\text{Nom}}_{l}}$ from the current–temperature fit. The two inequalities account for bidirectional flow, ensuring $\tau_{l,t}$ is an upper bound of the actual conductor temperature.
III Conductor health-aware unit commitment
Adopting DLR pushes line temperatures toward thermal limits, amplifying ETO risk in the presence of forecast errors. This section first characterizes the effect of correlation between WP and DLR forecast errors on ETO risk, motivating the explicit consideration of correlation in risk-averse DLR models. Then we propose the CHA-UC framework, which minimizes the sum of expected conductor depreciation and system operating costs to balance reliability and economics. The CHA-UC timeline is shown in Figure 3.
III-A Correlation of DLR and WP Forecast Errors
This section characterizes how WP–DLR forecast error correlation affects ETO-induced conductor depreciation.
Since the depreciation cost is a convex function of thermal stress, we recall the following lemma from [Shaked2007].
Lemma 1.
Let $X$, $Y$ be two random variables with the same mean. If $Y$ is more dispersed than $X$ in a mean-preserving spread, then for any non-decreasing convex function $g(\cdot)$,
$\mathbb{E}[g(Y)]\geq\mathbb{E}[g(X)].$ (9)
Here, let the WP and DLR forecast errors be
$e_{w}$ $=\hat{P}^{\text{Wind}}_{w}-P^{\text{Wind}}_{w},\,\,\,e_{l}=\hat{\overline{F}}^{\text{DLR}}_{l}-\overline{F}^{\text{DLR}}_{l},$ (10)
and given $s_{l}=\mathrm{sign}(p_{l}),$ let the stochastic part of the thermal stress on line $l$ induced by wind farm $w$ be
$\tilde{S}_{l,w}=A_{l}\bigl(e_{l}-s_{l}\mathrm{GSF}_{w,l}e_{w}\bigr),$ (11)
which captures the impact of forecast errors on line loading.
Proposition 2.
For a wind farm–line pair $(w,l)$, if $\mathrm{GSF}_{w,l}p_{l}<0$, then increasing correlation $\rho_{w,l}$ between $e_{w}$ and $e_{l}$ increases the dispersion of $\tilde{S}_{l,w}$. If $\mathrm{GSF}_{w,l}p_{l}>0$, then increasing correlation $\rho_{w,l}$ decreases the dispersion of $\tilde{S}_{l,w}$.
Proof.
The proof is presented in Appendix -A
Combining Proposition 2 with Lemma 1, we obtain the following result on expected depreciation cost.
Corollary 3.
For fixed marginal distributions of $e_{w}$ and $e_{l}$, increasing positive correlation $\rho_{w,l}$ increases $\mathbb{E}[C_{l}^{\mathrm{Dep}}]$ if $\mathrm{GSF}_{w,l}p_{l}<0$, and decreases $\mathbb{E}[C_{l}^{\mathrm{Dep}}]$ if $\mathrm{GSF}_{w,l}p_{l}>0$.
Proof.
The proof is presented in Appendix -B
Thus, the correlation between WP and DLR forecast errors exerts a systematic influence on line depreciation costs, depending on network topology and line flow conditions. Wind is the dominant common driver of both WP and DLR, creating strong correlation between their forecast errors, especially for nearby assets. Therefore, this effect should be accounted for in DLR operation, and the implementation of DLR on wind-connected lines. Notably, for typical WP tie lines, $\mathrm{GSF}_{w,l}p_{l}>0$ holds, so the correlation tends to hedge ETO risk because the wind farm is the only directly connected source.
III-B Baseline UC model
Before presenting the CHA-UC framework, the deterministic baseline model for DLR and SLR is introduced. Day-ahead UC and real-time operation is reflected in the model. The UC problem fixes commitment decisions day before the actual operation, using day-ahead forecasts and considering physical constraints of generators. The enforcement of UC constraints lead to a more comprehensive benchmark of results [Park2018].
Figure 3: Schematic of the CHA-UC optimization timeline. Processes that require actions outside optimization is colored with gray.
@@ III-B Baseline UC model
which is the sum of production cost, startup cost and the shut down cost. $\Xi:=\{\mathbf{g,p,u,}\,\bm{\theta,\delta}\}$
The day-ahead UC constraints, hereafter referred to as the UC constraints, are adapted from [Morales2013]:
@@
$\begin{split}&\sum_{l\mid s(l)=b}\hskip-4.2679ptp_{l,t}-\hskip-8.53581pt\sum_{l\mid r(l)=b}\hskip-4.2679ptp_{l,t}+\sum_{i\in\mathcal{I}_{b}}g_{i,t}+\hskip-8.53581pt\sum_{w\in\mathcal{W}_{b}}(\hat{P}^{\rm wind}_{w,t}-p^{\text{cur}}_{w,t})\\ &=P^{\text{d}}_{b,t}-p^{\text{shed}}_{b,t},\quad\forall b\in\mathcal{B},t\in\mathcal{T},\end{split}$ (14b)
@@
Once the day-ahead dispatch is fixed, the real-time operation minimizes the additional cost of balancing actual wind generation and weather realizations:
@@
where $\Xi:=\{\mathbf{g}^{\mathrm{adj}},\mathbf{p}^{\mathrm{adj}},\mathbf{r}^{\mathrm{+,-}},\bm{\theta}^{\mathrm{adj}}\}$. Reserve activations are penalized through $C_{i}^{\mathrm{r}\pm}$, and the quadratic term captures nonlinear production costs. This objective represents the operator’s total adjustment cost in real time. The constraints of the operation stage are as follows:
@@
III-C Conductor health-aware unit commitment
Given the linear fit and the piecewise depreciation cost function in Section II, the objective function of the CHA-UC day-ahead problem is formulated:
$\min_{\Xi}\quad\sum_{t\in\mathcal{T}}\sum_{i\in\mathcal{I}}\bigg[C^{\text{P}}_{i}(g_{i,t})u_{i,t}+\sum_{o\in\mathcal{O}_{i}}C^{\text{SU}}_{i,o}\delta_{i,o,t}+C^{\text{SD}}_{i}u^{\text{SD}}_{i,t}\bigg]$
$+\sum_{t\in\mathcal{T}}\sum_{b\in\mathcal{B}}C^{\text{VoLL}}p^{\text{shed}}_{b,t}+\sum_{s\in\mathcal{S}}\mathrm{Pr}_{s}\sum_{t\in\mathcal{T}}\bigg[\underbrace{\sum_{l\in\mathcal{L}^{\text{DLR}}}C^{\text{Dep}}_{l,t}(\hat{\tau}_{l,t,s})}_{\mathbb{E}(\text{Depreciation cost})}$
$+\underbrace{\sum_{b\in\mathcal{B}}C^{\text{VoLL}}\hat{p}^{\text{shed,adj}}_{b,t,s}}_{\mathbb{E}(\text{Load shedding cost})}+\underbrace{\sum_{i\in\mathcal{I}}\left[C^{\text{r+}}\hat{r}^{\text{+}}_{i,t,s}+C^{\text{r-}}\hat{r}^{\text{-}}_{i,t,s}\right]}_{\mathbb{E}(\text{Reserve activation cost})}\bigg].$ (17)
The objective includes the operation cost from (12) and the stochastic costs: the expected conductor depreciation cost, load shed costs and reserve activation costs. Here, $C^{\text{Dep}}_{l,t}$ is implemented using SOS2 constraints. All terms including the depreciation cost in the objective functions uses monetary units (\$), which eliminates the need for any additional scaling or conversion coefficient, simplifying the optimization and improving interpretability. The decision variables are identical to the baseline. The sample average approximation was used to calculate the expected value.
@@ III-C Conductor health-aware unit commitment
$\begin{split}&\sum_{l\mid s(l)=b}\hskip-4.2679pt\hat{p}^{\text{adj}}_{l,t,s}-\sum_{l\mid r(l)=b}\hskip-4.2679pt\hat{p}^{\text{adj}}_{l,t,s}+\sum_{i\in\mathcal{I}_{b}}\hat{g}^{\text{adj}}_{i,t,s}+\hskip-4.2679pt\sum_{w\in\mathcal{W}_{b}}\hskip-4.2679pt(\hat{P}^{\rm wind}_{w,t}\\ &-\hat{p}^{\text{cur,adj}}_{w,t,s}-\omega_{w,t,s})=P^{\text{d}}_{b,t}-\hat{p}^{\text{shed,adj}}_{b,t,s},\quad\forall b\in\mathcal{B},t\in\mathcal{T},\end{split}\hskip-8.53581pt$ (19f)
@@
The surrogate of the “wait-and-see” operation stage is formed to reflect the stochasticity in the “here-and-now” day-ahead UC. The surrogate corresponding to (16a)–(16g) are (19a)–(19g), respectively. They all represent the same operational constraints, but the latter are not actual; they represent the constraints with scenario realizations. Random variables regarding each scenario, $\omega_{w,t,s}$ and $\xi_{l,t,s}$ represent the wind forecast error and DLR forecast error. Equation (19h) and (19i) imposes a lower bound of $\tau_{l,t,s}$ using the linear conductor-temperature approximation (8). $\tau_{l,t,s}$ will be fixed to the lower bound as $C^{\text{Dep}}_{l,t}$ is a non-decreasing convex function, working as a proxy of conductor temperature. This linear approximation offers two advantages: it represents DLR forecast errors as equivalent flow deviations and handles bidirectional line flow efficiently.
@@
Analogous to equations (19h)-(19i), equations (22a)-(22b) enforce $\tau_{l,t}$ to reflect the conductor temperature before the operation stage.
TABLE I: Summarized results of Case Study I and Case Study II updated from the CSV files. The best result for each cost is in bold.
@@
Spring CHA 14.29 0.08 0.22 11.30 14.60 14.68 1.89 1.35 0.66 96.8 18.58
DLR 14.16 1.23 7.6 15.39 13.73 2.71 3.89 11.72 66.2 32.05
SLR 14.57 0.13 92.7 14.70 14.15 2.28 2.83 8.80 85.9 28.06
QRF 14.25 0.72 29.2 14.97 13.81 2.63 3.75 8.67 74.1 28.86
Summer CHA 29.96 0.12 0.12 14.3 30.21 29.77 2.92 2.37 0.28 54.3 35.34
DLR 29.63 0.97 9.8 30.60 28.48 4.11 5.44 10.96 28.9 49.00
SLR 29.96 0.46 56.5 30.42 28.84 3.67 4.29 8.46 34.6 45.26
QRF 30.35 0.19 24.1 30.55 29.17 3.93 4.89 7.44 28.1 45.43
Fall CHA 22.08 0.06 0.01 26.7 22.15 21.99 2.56 1.30 0.16 83.8 26.01
DLR 21.85 0.48 24.8 22.33 20.89 3.66 3.48 12.62 46.9 40.65
SLR 22.26 0.03 126.7 22.30 21.28 3.26 2.80 7.32 72.9 34.66
QRF 22.01 0.07 47.6 22.08 21.04 3.57 3.38 8.60 38.9 36.58
Winter CHA 18.50 0.05 0.13 6.7 18.68 18.63 2.06 0.70 0.14 51.5 21.53
DLR 18.39 0.67 4.4 19.06 17.72 3.00 3.01 8.94 41.7 32.67
SLR 18.86 0.00 53.5 18.86 18.15 2.62 2.02 5.30 59.8 28.08
QRF 18.44 0.30 4.4 18.74 17.75 2.98 2.76 7.46 44.5 30.95
Annual CHA 21.23 0.08 0.12 14.8 21.44 21.29 2.36 1.44 0.31 71.8 25.41
DLR 21.04 0.84 11.7 21.87 20.23 3.38 3.97 11.08 46.0 38.65
SLR 21.44 0.16 82.6 21.60 20.63 2.96 3.00 7.49 63.3 34.08
QRF 21.29 0.32 26.5 21.61 20.47 3.28 3.71 8.04 46.4 35.50
IV Case Study
Two case studies are performed on the synthetic Texas 123-bus backbone transmission system (TX-123BT) [Lu2025] using 2021-2022 weather data. Case Study I runs an annual simulation with uncertain DLR forecasts (with errors) and perfect wind power (WP) forecasts (no errors) to illustrate CHA-UC. Case Study II evaluates CHA-UC in a more realistic setting with uncertainty in both WP and DLR forecasts to analyze their correlation effects.
((a))
((b))